# The CLAS Collaboration

### v2

@@ III.5.4 $\pi^{-}$ missing topology
where $P_{i}^{\mu}$ are the four-momenta of particle $i$ and $X$ corresponds to the undetected part.
To isolate quasi-free events in the $\pi^{-}$ missing topology, a special procedure was developed, in which the following quantity was used to perform the exclusivity cut,
@@
Figure 7 shows the distribution of the quantity $M_{X[\pi^{-}]}$ plotted for the experimental data (black histogram) and Monte Carlo simulation (blue histogram) in one 25-MeV-wide $W$ bin. The magenta histogram shows the difference between them and thus represents the distribution of background events, which are mainly events affected by FSIs with the spectator. The green line corresponds to the cut applied to select quasi-free events.
Figure 7: Distributions of the quantity $M_{X[\pi^{-}]}$ (defined by Eq. (4)) in one 25-MeV-wide $W$ bin for the experimental data (black histogram), Monte Carlo simulation (blue histogram), and their difference (magenta histogram). The explanation of the fit curves is given in the text. The green line shows the applied exclusivity cut.
As seen in Fig. 7, the exclusivity cut does not allow for complete isolation of the quasi-free event sample. Tightening the cut would lead to significant reduction in the statistics of selected events, yet without total elimination of the FSI-background. Therefore, an “effective correction” of the FSI-background admixture was performed, which included the following steps.
  • The $M_{X[\pi^{-}]}$ distribution of the reconstructed Monte Carlo events (blue histogram) was fit with a polynomial. A typical result of this fit is shown in Fig. 7 by the solid orange curve.
  • The magenta background distribution was fit with a Gaussian. The result of the fit is shown by the dark-magenta dash-dotted curve.
  • The orange and dark-magenta curves were summed up to produce the red dashed curve that matches the black experimental histogram.
@@
  • In each $W$ bin, the experimental event yield was multiplied by the factor $F_{\text{\tiny FSI}}$, which served as an effective correction due to the remaining admixture of the FSI-background events.
The factor $F_{\text{\tiny FSI}}$ is assumed to be only $W$ dependent as it was not found to exhibit any $Q^{2}$ dependence, and the dependence on the final hadronic variables is neglected due to the statistics limitation. The value of $F_{\text{\tiny FSI}}$ varies from $\sim$0.97 to $\sim$0.93 for the $W$ bins in the range from 1.45 GeV to 1.825 GeV. For $W<1.45$ GeV, the correction is not needed as no mismatch between the experimental and simulated distributions is observed in this region (see Sec. VII.3).
Note that the exclusivity cut shown in Fig. 7, accompanied by the corresponding correction, accounts for all other possible effects that along with the FSI effects may contribute to the mismatch between the data and the simulation in this topology (including the minor three-pion background contribution).
@@ IV.1 Fermi smearing of the invariant mass $W$
$W_{\text{i}}$ $=$ $\sqrt{(P_{p}^{\mu}+P_{\gamma_{v}}^{\mu})^{2}}$ (6)
$W_{\text{f}}$ $=$ $\sqrt{(P_{\pi^{+}}^{\mu}+P_{\pi^{-}}^{\mu}+P_{p^{\prime}}^{\mu})^{2}}$ (7)
Here $P_{\pi^{+}}^{\mu}$, $P_{\pi^{-}}^{\mu}$, and $P_{p^{\prime}}^{\mu}$ are the four-momenta of the final-state hadrons, $P_{p}^{\mu}$ is the four-momentum of the initial proton and $P_{\gamma_{v}}^{\mu}=P_{e}^{\mu}-P_{e^{\prime}}^{\mu}$ the four-momentum of the virtual photon with $P_{e}^{\mu}$ and $P_{e^{\prime}}^{\mu}$ the four-momenta of the incoming and scattered electrons, respectively.
In general, to determine $W_{\text{f}}$, all final hadrons should be registered, while for the calculation of $W_{\text{i}}$, it is sufficient to register just the scattered electron. The latter option allows one to analyze event samples in which information on the reaction final state is incomplete, as e.g. in topologies with one unregistered final hadron (see Sec. III.5.1). However, in reactions off bound nucleons, this opportunity comes with a complication as to correctly calculate $W_{\text{i}}$, information on the initial proton momentum ($P_{p}^{\mu}$) is also required. In this analysis, however, this information is not accessible in the $\pi^{-}$ missing topology. This situation brings up the choice to either demand registration of all final hadrons to determine $W_{\text{f}}$ (which reduces the analysis flexibility) or to calculate $W_{\text{i}}$ assuming the initial proton to be at rest.
In this analysis, the invariant mass $W_{\text{i}}$ calculated under the target-at-rest assumption was used to describe the reaction, based on the statistically dominant $\pi^{-}$ missing topology. For this reason, the extracted cross sections turn out to be Fermi smeared Skorodumina et al. (2015b, 2017a). To retrieve the non-smeared observable, a correction that unfolds this effect should be applied, which is described in Sec. V.3.
IV.2 Lab to CMS transformation
Once the double-pion events were selected as described in Section III, the laboratory four-momenta of all final particles are known as they are either registered or reconstructed as missing. These four-momenta were then used to calculate the kinematic variables, which are introduced in Sec. IV.3.
@@ IV.2 Lab to CMS transformation
The CMS is uniquely defined as the system where the initial proton and the photon move towards each other with the $z_{CMS}$-axis pointing along the photon and the net momentum equal to zero. However, the procedure of the Lab to CMS transformation differs depending on the specificity of the reaction initial state (real or virtual photons, at rest or moving initial proton).
The correct procedure of the Lab to CMS transformation for an electroproduction experiment off a moving proton can be subdivided into two major steps.
@@
    First, one needs to perform a transition to the auxiliary system, where the target proton is at rest, while the incoming electron moves along the $z$-axis. This system can be called “quasi-Lab”, since the initial conditions of the reaction in this frame imitate those existing in the Lab system in the case of the free proton experiment. The recipe of the Lab to quasi-Lab transformations is given in detail in Ref. Skorodumina et al. (2017a).
@@
    Then, the quasi-Lab to CMS transformation should be performed by the standard method used for an electroproduction experiment off a proton at rest Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina .
The first step of this procedure (Lab to quasi-Lab transformation) implies that the momentum of the initial proton is known for each reaction event Skorodumina et al. (2017a). In this analysis, however, information on the initial proton momentum can be accessible only in the fully exclusive topology (it can be deduced via momentum conservation as shown in Sec. III.5.3), while in the $\pi^{-}$ missing topology, this information turns out to be irrevocably lost due to incomplete knowledge about the reaction final state. As a result, for the majority of the analyzed events, the correct Lab to CMS transformation could not be performed. For this reason, in this analysis, the procedure of Lab to CMS transformation for an electroproduction experiment off a proton at rest was used. The procedure is described in Refs. Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina and was employed for both fully exclusive and $\pi^{-}$ missing topologies for consistency.
This approximation in the Lab to CMS transformation introduces a systematic inaccuracy to the extracted cross sections. A correction for this effect is included in the procedure of unfolding the effects of the initial proton motion (see Sec. V.3).
@@ IV.3 Kinematic variables
Meanwhile, the three-body final hadron state of the reaction is unambiguously determined by five kinematic variables Fedotov et al. (2017), and in general there can be different options for their choice. In this analysis, the following generalized set of variables was used Fedotov et al. (2017, 2018); Byckling and Kajantie (1971); Isupov et al. (2017); Fedotov et al. (2009); Mokeev et al. (2016); Skorodumina et al. (2020); Skorodumina :
@@
  • the angle $\alpha_{h_{1}}$ between the two planes: (i) defined by the three-momenta of the virtual photon (or initial proton) and the first final hadron and (ii) defined by the three-momenta of all final hadrons.
In this study, the cross sections were obtained in three sets of variables depending on various assignments for the first, second, and third final hadrons:
@@
$M_{h_{1}h_{2}}$
Invariant mass
8 10 12 12
$M_{h_{2}h_{3}}$
Invariant mass
8 10 12 12
$\theta_{h_{1}}$
Polar angle
6 8 10 10
$\varphi_{h_{1}}$
Azimuthal angle
5 5 5 6
$\alpha_{h_{1}}$
Angle between planes
5 6 8 8
@@ IV.4 Binning and kinematic coverage
[…]
The kinematic coverage in the final state variables has the following reaction-related features. The angular variables $\theta_{h_{1}}$, $\varphi_{h_{1}}$, and $\alpha_{h_{1}}$ vary in the fixed ranges of $[0,~{}\pi]$, $[0,~{}2\pi]$, and $[0,~{}2\pi]$, respectively. Meanwhile, the ranges of the invariant masses $M_{h_{1}h_{2}}$ and $M_{h_{2}h_{3}}$ are not fixed -- they are $W$ dependent and broaden as $W$ grows.
@@
Figure 8: $Q^{2}$ versus $W$ distribution populated with the selected double-pion events. The cross section was calculated in 2D cells within the white boundaries.
@@
$M_{\text{lower}}$ $=m_{h_{1}}+m_{h_{2}}$ (8)
$M_{\text{upper}}(W)$ $=W-m_{h_{3}}$
@@
Since the cross section is calculated in $W$ bins of a given width, the boundary of $M_{\text{upper}}$ is not distinct. For the purpose of binning in mass, the value of $M_{\text{upper}}$ was calculated using $W_{\text{center}}$, at the center of the $W$ bin, which caused events with $W>W_{\text{center}}$ to be located beyond $M_{\text{upper}}$. For this reason, it was decided to use a specific arrangement of mass bins with the bin width $\Delta M$ determined by
@@
The chosen arrangement of bins forces the last bin to be situated completely out of the boundaries calculated according to Eq. (8) using $W_{\text{center}}$. The cross section in this extra bin is not reported. However, this bin was kept in the analysis since its contents (though being very small) contribute to all cross sections that were obtained by integrating over the corresponding invariant mass distribution.
Note that the cross section in the next to last bin in invariant mass needs a correction. See more details in Sec. V.4.
@@ IV.5.1 Electron scattering cross section
  • $N_{\text{full}}$ and $N_{\text{empty}}$ are the numbers of selected double-pion events inside the seven-dimensional bin for runs with liquid deuterium and empty target, respectively;
@@ IV.5.2 Virtual photoproduction cross section
The limited statistics of the experiment do not allow for estimates of the five-fold differential cross section $\sigma_{\text{v}}$ with reasonable accuracy. Therefore, the cross section $\sigma_{\text{v}}$ was first obtained on the multi-dimensional grid and then integrated over at least four hadronic variables, which means that only single-differential and fully integrated cross sections were obtained.
For each $W$ and $Q^{2}$ bin, the following cross sections were extracted for each set of variables (see Sec. IV.3),
@@
Regarding the middle row in Eq. (15), note the following. Although being differential in $[-\cos\theta]$, the cross sections were calculated in $\Delta\theta$ bins, which are of equal size in the corresponding $W$ subrange (see also Sec. IV.4). This follows the convention used to extract the $\theta$-distributions in studies of double-pion cross sections Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018).
@@ IV.6 Efficiency evaluation
In this study, the Monte Carlo simulation was performed using the TWOPEG-D event generator Skorodumina et al. (2017a), which is capable of simulating the quasi-free process of double-pion electroproduction off a moving proton. In this event generator, the Fermi motion of the initial proton is generated according to the Bonn potential Machleidt et al. (1987) and then is naturally merged into the specific kinematics of double-pion electroproduction. TWOPEG-D accounts for radiative effects according to the approach described in Refs. Mo and Tsai (1969); Skorodumina et al. (2017b).
Events generated with TWOPEG-D were passed through the standard detector simulation and reconstruction procedures with the majority of parameters kept the same as in the studies Fedotov et al. (2017, 2018); Markov et al. (2014); Ye Tian and Gothe (2020); Ye Tian et al. (2023), which were also devoted to the “e1e” run period. More information on the simulation/reconstruction procedure and the related parameters can be found in Ref. Skorodumina et al. (2020).
[…]
TWOPEG-D performs a weighted event generation Skorodumina et al. (2017b), which means that all kinematic variables are randomly generated according to the double-pion production phase space, while each event generated at a particular kinematic point acquires an individual weight that reflects the cross section value at that point. The efficiency factor $\mathcal{E}$ from Eq. (10) was then calculated in each $\Delta W\Delta Q^{2}\Delta^{5}\tau$ bin by
@@
[…]
In some kinematic bins, the efficiency $\mathcal{E}$ could not be reliably determined due to boundary effects, bin-to-bin event migration, and/or limited Monte Carlo statistics. In such bins, the relative efficiency uncertainty $\delta\mathcal{E}/\mathcal{E}$ is typically large. In this study, a cut on the relative efficiency uncertainty $\delta\mathcal{E}/\mathcal{E}$ was performed that excluded from consideration all multi-dimensional cells with uncertainties greater than 30%. The excluded cells were ranked as “empty cells” and, along with other empty cells, were subject to the filling procedure (see Sec. V.1).
The cut on the relative efficiency uncertainty directly impacts the cross section uncertainties. On the one hand, it eliminates the $\Delta^{5}\tau$ bins with large $\delta\mathcal{E}/\mathcal{E}$ values, thus reducing the total statistical uncertainty of the extracted cross sections (see Sec. VI.1). On the other hand, this cut increases the number of empty cells, thus increasing the cross section model dependence and the uncertainty associated with it (see Sec. VI.2). The cut value was chosen as a compromise between these two effects.
The idea of this cut has been taken from the free-proton study Fedotov et al. (2017, 2018), which also sets the cut value at 30%. More details can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).

### v4

@@ III.5.4 $\pi^{-}$ missing topology
where $P_{i}^{\mu}$ are the four-momenta of the particle $i$ and $X$ corresponds to the undetected part.
To isolate quasifree double-pion events in the $\pi^{-}$ missing topology, a special procedure was developed, in which the following quantity was used to perform the exclusivity cut:
@@
Figure 7 shows the distribution of the quantity $M_{X[\pi^{-}]}$ plotted for the experimental data (black histogram) and the Monte Carlo simulation (blue histogram) in one 25-MeV-wide $W$ bin. The magenta histogram shows the difference between them and thus represents the distribution of background events, which are mainly events affected by FSIs with the spectator. The green line corresponds to the cut applied to select quasifree events.
Figure 7: Distributions of the quantity $M_{X[\pi^{-}]}$ [defined by Eq. (4)] in one 25-MeV-wide $W$ bin for the experimental data (black histogram), the Monte Carlo simulation (blue histogram), and their difference (magenta histogram). The black and blue histograms are normalized to their maxima. The explanation of the fit curves is given in the text. The vertical green line shows the applied exclusivity cut.
As seen in Fig. 7, the exclusivity cut does not allow for complete isolation of the quasifree event sample. Tightening the cut would lead to significant reduction in the statistics of selected events, yet without total elimination of the FSI background. Therefore, an “effective correction” of the FSI-background admixture was performed, which included the following steps.
  • The $M_{X[\pi^{-}]}$ distribution of reconstructed Monte Carlo events (blue histogram) was fit with a polynomial. A typical result of this fit is shown in Fig. 7 by the solid orange curve.
  • The magenta background distribution was fit with a Gaussian. The result of the fit is shown by the dark magenta dash-dotted curve.
  • The orange and dark magenta curves were summed up to produce the red dashed curve that matches the black experimental histogram.
@@
  • In each $W$ bin, the experimental event yield was multiplied by the factor $F_{\text{\tiny FSI}}$, which served as an effective correction due to the remaining admixture of FSI-background events.
The factor $F_{\text{\tiny FSI}}$ is assumed to be only $W$ dependent as it was not found to exhibit any $Q^{2}$ dependence, and the dependence on the final hadronic variables is neglected due to the statistics limitation. The value of $F_{\text{\tiny FSI}}$ varies from $\approx$0.97 to $\approx$0.93 for the $W$ bins in the range from 1.45 to 1.825 GeV. For $W<1.45$ GeV, the correction was not needed as no mismatch between the experimental and the simulated distributions was observed in this region (see Sec. VII.3).
Note that the exclusivity cut shown in Fig. 7, accompanied by the corresponding correction, accounts for all other possible effects that along with FSI effects may contribute to the mismatch between the data and the simulation in this topology (including a minor three-pion background contribution).
@@ IV.1 Fermi smearing of the invariant mass $W$
$W_{\text{i}}$ $=$ $\sqrt{(P_{p}^{\mu}+P_{\gamma_{v}}^{\mu})^{2}}.$ (6)
$W_{\text{f}}$ $=$ $\sqrt{(P_{\pi^{+}}^{\mu}+P_{\pi^{-}}^{\mu}+P_{p^{\prime}}^{\mu})^{2}}.$ (7)
Here $P_{\pi^{+}}^{\mu}$, $P_{\pi^{-}}^{\mu}$, and $P_{p^{\prime}}^{\mu}$ are the four-momenta of the final-state hadrons, $P_{p}^{\mu}$ is the four-momentum of the initial proton, and $P_{\gamma_{v}}^{\mu}=P_{e}^{\mu}-P_{e^{\prime}}^{\mu}$ the four-momentum of the virtual photon with $P_{e}^{\mu}$ and $P_{e^{\prime}}^{\mu}$ the four-momenta of the incoming and scattered electrons, respectively.
In general, to determine $W_{\text{f}}$, all final hadrons should be registered, while for calculation of $W_{\text{i}}$, it is sufficient to register just the scattered electron. The latter option allows one to analyze event samples in which information on the reaction final state is incomplete, as, e.g. in topologies with one unregistered final hadron (see Sec. III.5.1). However, in reactions off bound nucleons, this opportunity comes with a complication, for to correctly calculate $W_{\text{i}}$, information on the initial proton momentum ($P_{p}^{\mu}$) is also required. In this analysis, however, this information is not accessible in the $\pi^{-}$ missing topology. This situation brings up a choice to either demand registration of all final hadrons to determine $W_{\text{f}}$ (which reduces the analysis flexibility) or to calculate $W_{\text{i}}$ assuming the initial proton to be at rest.
In this study, the invariant mass $W_{\text{i}}$ calculated under the target-at-rest assumption was used to describe the reaction, based on the statistically dominant $\pi^{-}$ missing topology. For this reason, the extracted cross sections turn out to be Fermi smeared Skorodumina et al. (2015b, 2017a). To retrieve the nonsmeared observable, a correction that unfolds this effect was applied, which is described in Sec. V.3.
IV.2 Lab-to-CMS transformation
Once the double-pion events were selected as described in Sec. III, the laboratory four-momenta of all final particles are known as they are either registered or reconstructed as missing. These four-momenta were then used to calculate the kinematic variables, which are introduced in Sec. IV.3.
@@ IV.2 Lab-to-CMS transformation
The CMS is uniquely defined as the system where the initial proton and the photon move towards each other with the net momentum equal to zero and the $z_{CMS}$ axis pointing along the photon. However, the procedure of the Lab-to-CMS transformation differs depending on the specificity of the reaction initial state (real or virtual photons, at rest or moving initial proton).
The correct procedure of the Lab-to-CMS transformation for an electroproduction experiment off a moving proton can be subdivided into two major steps.
@@
    First, one needs to perform a transition to the auxiliary system, where the target proton is at rest, while the incoming electron moves along the $z$ axis. This system can be called “quasiLab” because the initial conditions in this frame imitate those existing in the Lab system for the case of a free proton experiment. The recipe of the Lab-to-quasiLab transformation is given in detail in Ref. Skorodumina et al. (2017a).
@@
    Then, the quasiLab-to-CMS transformation should be performed by the standard method used for an electroproduction experiment off a proton at rest Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina .
The first step of this procedure (Lab-to-quasiLab transformation) implies that the momentum of the initial proton is known for each analyzed event Skorodumina et al. (2017a). In this study, however, information on the initial proton momentum can be accessible only in the fully exclusive topology (it can be deduced via momentum conservation as shown in Sec. III.5.3), while in the $\pi^{-}$ missing topology, this information turns out to be irrevocably lost due to incomplete knowledge about the reaction final state. As a result, for the majority of analyzed events, the correct Lab-to-CMS transformation could not be performed. For this reason, in this analysis, the procedure of the Lab-to-CMS transformation for an electroproduction experiment off a proton at rest was used. The procedure is described in Refs. Fedotov et al. (2017, 2018); Skorodumina et al. (2020); Skorodumina and was employed for both the fully exclusive and the $\pi^{-}$ missing topologies for consistency.
This approximation in the Lab-to-CMS transformation introduces a systematic inaccuracy to the extracted cross sections. A correction for this effect is included in the procedure of unfolding the effects of the initial proton motion (see Sec. V.3).
@@ IV.3 Kinematic variables
Meanwhile, the three-body final hadronic state of the reaction is unambiguously determined by five kinematic variables Fedotov et al. (2017), and in general there can be different options for their choice. In this analysis, the following generalized set of variables was used Fedotov et al. (2017, 2018); Byckling and Kajantie ; Isupov et al. (2017); Fedotov et al. (2009); Mokeev et al. (2016); Skorodumina et al. (2020); Skorodumina :
@@
  • the angle $\alpha_{h_{1}}$ between the two planes: (a) defined by the three-momenta of the virtual photon (or initial proton) and the first final hadron and (b) defined by the three-momenta of all final hadrons.
The cross sections were obtained in three sets of variables depending on various assignments for the first, second, and third final hadrons:
@@
$M_{h_{1}h_{2}}$ Invariant mass 8 10 12 12
$M_{h_{2}h_{3}}$ Invariant mass 8 10 12 12
$\theta_{h_{1}}$ Polar angle 6 8 10 10
$\varphi_{h_{1}}$ Azimuthal angle 5 5 5 6
$\alpha_{h_{1}}$ Angle between planes 5 6 8 8
@@ IV.4 Binning and kinematic coverage
The available kinematic coverage in the initial-state variables is shown by the $Q^{2}$ versus $W$ distribution in Fig. 8. This distribution is filled with the double-pion events that survived after the event selection described in Sec. III. The white boundary shows the analyzed kinematic area, where the double-pion cross sections were extracted. The black grid demonstrates the chosen binning in the initial-state variables (25 MeV in $W$ and 0.05 GeV2 in $Q^{2}$).
The kinematic coverage in the final-state variables has the following reaction-related features. The angular variables $\theta_{h_{1}}$, $\varphi_{h_{1}}$, and $\alpha_{h_{1}}$ vary in the fixed ranges of $[0,~{}\pi]$, $[0,~{}2\pi]$, and $[0,~{}2\pi]$, respectively. Meanwhile, the ranges of the invariant masses $M_{h_{1}h_{2}}$ and $M_{h_{2}h_{3}}$ are not fixed — they are $W$ dependent and broaden as $W$ grows Skorodumina et al. (2020); Skorodumina .
@@
Figure 8: $Q^{2}$ versus $W$ distribution populated with selected double-pion events. The black grid demonstrates the chosen binning in the initial-state variables (25 MeV in $W$ and 0.05 GeV2 in $Q^{2}$) The cross section was calculated in two-dimensional cells within the white boundaries.
@@
$M_{\text{lower}}$ $=m_{h_{1}}+m_{h_{2}};$ (8)
$M_{\text{upper}}(W)$ $=W-m_{h_{3}}.$
@@
Since the cross section is calculated in $W$ bins of a given width, the boundary of $M_{\text{upper}}$ is not distinct. For the purpose of binning in masses, the value of $M_{\text{upper}}$ was calculated using $W_{\text{center}}$, at the center of the $W$ bin, which caused events with $W>W_{\text{center}}$ to be located beyond $M_{\text{upper}}$. For this reason, it was decided to use a specific arrangement of mass bins with the bin width $\Delta M$ determined by
@@
The chosen arrangement of bins forces the last mass bin to be situated completely out of the boundaries determined according to Eq. (8) using $W_{\text{center}}$. The cross section in this extra bin is not reported. However, the bin was kept when integrating the cross section over the mass distributions so that no events were lost [see Eqs. (14)].
Note that the cross section in the next-to-last bin in invariant masses needs a correction. See more details in Sec. V.4.
@@ IV.5.1 Electron scattering cross section
  • $N_{\text{full}}$ and $N_{\text{empty}}$ are the numbers of selected double-pion events inside a seven-dimensional bin for runs with liquid deuterium and empty target, respectively;
@@ IV.5.2 Virtual photoproduction cross section
The limited statistics of the experiment did not allow for estimates of the five-fold differential cross section $\sigma_{\text{v}}$ with reasonable accuracy. Therefore, the cross section $\sigma_{\text{v}}$ was first calculated on the multi-dimensional grid and then integrated over at least four hadronic variables, which means that only single-differential and fully integrated cross sections were obtained.
For each variable set described in Sec. IV.3, the following cross sections were extracted:
@@
Regarding the middle row in Eq. (15), note the following. Although being differential in $[-\cos\theta]$, the cross sections were calculated in $\Delta\theta$ bins, which are of equal size in a corresponding $W$ subrange (see also Sec. IV.4). This follows the convention used to extract the $\theta$ distributions in studies of double-pion cross sections Ripani et al. (2002, 2003); Fedotov et al. (2007, 2009); Isupov et al. (2017); Trivedi and Gothe (2018); Fedotov et al. (2017, 2018).
@@ IV.6 Efficiency evaluation
In this study, the Monte Carlo simulation was performed using the TWOPEG-D event generator Skorodumina et al. (2017a), which is capable of simulating a quasifree process of double-pion electroproduction off a moving proton. In this event generator, the Fermi motion of the initial proton is generated according to the Bonn potential Machleidt et al. (1987) and then is naturally merged into the specific kinematics of double-pion electroproduction. TWOPEG-D accounts for radiative effects according to the approach described in Refs. Mo and Tsai (1969); Skorodumina et al. (2017b).
Events generated with TWOPEG-D were passed through a standard multistage procedure of the detector simulation and event reconstruction with the majority of the parameters kept the same as in the studies Fedotov et al. (2017, 2018); Markov et al. (2014); Ye Tian and Gothe (2020); Ye Tian et al. (2023) that were also devoted to the “e1e” run period. More information on the procedure and the related parameters can be found in Ref. Skorodumina et al. (2020).
In studies of double-pion cross sections, it is important to generate enough Monte Carlo statistics in order to saturate each multi-dimensional bin of the reaction phase space with events (see Table 1). Insufficient Monte Carlo statistics will lead to an improper efficiency evaluation and an unnecessary rise in the number of empty cells (see Sec. V.1), thus systematically affecting the accuracy of the extracted cross sections. For this study, the total of about 4$\cdot$1010 double-pion events were generated in the investigated kinematic region, which was found to be sufficient.
TWOPEG-D performs a weighted event generation Skorodumina et al. (2017b). This means that all kinematic variables are randomly generated according to the double-pion production phase space, while each event generated at a particular kinematic point acquires an individual weight that reflects the cross section value at that point. The efficiency factor $\mathcal{E}$ from Eq. (10) was then calculated in each $\Delta W\Delta Q^{2}\Delta^{5}\tau$ bin by
@@
Figure 9: Extracted single-differential cross sections for the cases when the contribution from the empty cells was ignored (open squares) and when it was taken into account (black circles). The former are reported with the uncertainty $\delta_{\text{stat}}^{\text{tot}}$ given by Eq. (22) and the latter – with the uncertainty $\delta_{\text{stat,mod}}^{\text{tot}}$ given by Eq. (26). All distributions are given for one particular bin in $W$ and $Q^{2}$ (i.e. $W=$1.6375 GeV, $Q^{2}=$0.625 GeV2).
In some kinematic bins, the efficiency $\mathcal{E}$ could not be reliably determined due to boundary effects, bin-to-bin event migration, and/or limited Monte Carlo statistics. In such bins, the relative efficiency uncertainty $\delta\mathcal{E}/\mathcal{E}$ is typically large. In this study, a cut on the relative efficiency uncertainty $\delta\mathcal{E}/\mathcal{E}$ was performed that excluded from consideration all multi-dimensional cells with uncertainties greater than 30%. The excluded cells were ranked as “empty cells” and, along with the other empty cells, were subject to the filling procedure (see Sec. V.1).
It is noteworthy that the impact of the $\delta\mathcal{E}/\mathcal{E}$ cut on the cross section uncertainties is not straightforward. On the one hand, the cut eliminates $\Delta^{5}\tau$ bins with large $\delta\mathcal{E}/\mathcal{E}$ values, hence reducing the uncertainty due to the Monte Carlo statistics and, consequently, the total statistical uncertainty of the extracted cross sections (see Sec. VI.1). On the other hand, this cut increases the number of empty cells, thus increasing the cross section model dependence and the uncertainty associated with it (see Sec. VI.2). The cut value was chosen as a compromise between these two effects.
The idea of this cut was taken from the free-proton study Fedotov et al. (2017, 2018), which also sets the cut value at 30%. More details can be found in Refs. Skorodumina et al. (2020); Skorodumina ; Fedotov et al. (2017, 2018).